# SINDI: An Efficient Index for Sparse Vector Approximate Maximum Inner Product Search

### v2

@@ 3.2. Efficient Distance Computation
Figure 4 illustrates the search procedure. Since $\lVert\mathcal{D}\rVert=9$, the distance array $A$ is initialized with $\text{size}(A)=9$ and all elements set to 0. The query $\vec{q}$ contains three non-zero components, $q^{1}$, $q^{5}$, and $q^{8}$, so only the inverted lists $I_{1}$, $I_{5}$, and $I_{8}$ are traversed. Consider $\vec{x}_{4}$ as an example. From $I_{1}$, we obtain $x_{4}^{1}=6.8$, and the product $q^{1}\times x_{4}^{1}=17.0$ (this multiplication can be SIMD-accelerated) is temporarily stored in $T[0]$. Accumulating $T[0]$ into $A[4]$ gives $A[4]=0+17.0=17.0$. Similarly, from $I_{5}$ we compute $x_{4}^{5}\times q^{5}=14.0$, store it in $T[2]$, and add it to obtain $A[4]=31.0$. The computation for $I_{8}$ proceeds analogously. Finally, $A[4]$ becomes $36.1$, which equals $\delta(\vec{x}_{4},\vec{q})$. Although $\vec{x}_{4}$ has another non-zero entry $x_{4}^{2}$, it is not in $\Omega(\vec{x}_{4},\vec{q})$ and thus does not contribute to the inner product. The same accumulation process applies to $\vec{x}_{2}$, $\vec{x}_{3}$, $\vec{x}_{6}$, and $\vec{x}_{7}$. Eventually, $A[4]$ holds the largest value, so the nearest neighbor of $\vec{q}$ is $\vec{x}_{4}$.
Figure 4. An Example of Sindi and Query Process.
Using SIMD instructions, Sindi processes the $j$-th inverted list $I_{j}$ in batches, multiplying $q^{j}$ with each $x_{i}^{j}$ it contains and writing the results sequentially into $T^{j}$. This approach maximizes CPU utilization and reduces the complexity of the inner product computation from $\mathcal{O}(\lVert q\rVert+\lVert x\rVert)$ to $\mathcal{O}\!\left(\frac{\lVert q\rVert}{s}\right)$, where $s$ denotes the number of elements processed per SIMD operation. The complexity of Sindi’s distance computation is derived as follows:
Theorem 3.3 (Amortized Time Complexity of Sindi Distance Computation).
@@
Proof. The total number of non-zero entries accessed in all posting lists for $\mathcal{J}$ is
$\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert.$
Since $s$ entries can be processed in parallel using SIMD, the total time is
$T_{\text{total}}=\sum_{j\in\mathcal{J}}\frac{\lVert\mathcal{I}^{j}\rVert}{s}.$
Amortizing over all $\lVert\mathcal{X}\rVert$ candidates gives
$T=\frac{T_{\text{total}}}{\lVert\mathcal{X}\rVert}=\frac{\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert}{s\cdot\lVert\mathcal{X}\rVert}.$
For any $\vec{x}_{i}\in\mathcal{X}$, we have $q\cap x_{i}\subseteq q$, implying that $\|\Omega(\vec{x}_{i},\vec{q})\|\leq\|q\|$. Therefore:
$T\leq\frac{\sum_{\vec{x}_{i}\in\mathcal{X}}\|q\|}{s\cdot\|\mathcal{X}\|}=\frac{\|q\|\cdot\|\mathcal{X}\|}{s\cdot\|\mathcal{X}\|}=\frac{\|q\|}{s}.$
Substituting into $T$ yields
$T\leq\frac{\sum_{\vec{x}_{i}\in\mathcal{X}}\lVert q\rVert}{s\cdot\lVert\mathcal{X}\rVert}=\frac{\lVert q\rVert}{s}.$
Hence, the amortized per-vector complexity is
$\Theta\!\left(\frac{\lVert q\rVert}{s}\right).$
$\square$
In summary, the Sindi index removes the overhead of redundant dimension matching and exploits SIMD parallelism to compute inner products in batches, thereby fully utilizing CPU computational resources. It also eliminates the memory access costs of traversing the original vectors, as the required values are retrieved directly from the posting lists and used for in-place product computation during retrieval.
3.3. Cache Optimization
@@ 3.3.1. Window Switch
(1) Inner product computation. For each scanned posting list $I_{j,w}$, compute the products $q^{j}\times x_{i}^{j}$ for all its entries, using SIMD instructions when possible. These products are written sequentially into a temporary array $T^{j}$, which has the same length as $I_{j,w}$ and is index-aligned with it — i.e., $T^{j}[t]$ stores the product for the $t$-th posting in $I_{j,w}$. The accumulation stage then adds each $T^{j}[t]$ into the corresponding $A[i\bmod\lambda]$ using $\mathcal{O}(1)$ time per update.
@@
Note that Window Switch only changes the order of list entries scanned, without altering the number of arithmetic operations. Thus, the time complexity of distance computation remains $\mathcal{O}\!\left(\frac{\lVert q\rVert}{s}\right)$.
@@ 4    $w\leftarrow\lfloor\frac{i}{\lambda}\rfloor$
3.3.2. Construction and Search
The construction process of the full-precision Sindi index with Window Switch is shown in Algorithm 1. Given a sparse vector dataset $\mathcal{D}$ of dimension $d$, the algorithm iterates over each dimension $j$ (Line 1). For each $j$, it collects all vectors $\vec{x}_{i}\in\mathcal{D}$ having a non-zero entry $x_{i}^{j}$ into a temporary set $\mathcal{X}$ (Line 2). These vectors are then appended to the corresponding window $I_{j,w}$ of $I_{j}$ based on their IDs, where $w=\lfloor i/\lambda\rfloor$ (Lines 3–5). After processing all dimensions, the inverted index $I$ is returned (Line 6). The time complexity of the construction process is $\mathcal{O}(\lVert\mathcal{D}\rVert\lVert\bar{x}\rVert)$, where $\lVert\bar{x}\rVert=\frac{\sum_{\vec{x}_{i}\in\mathcal{D}}\lVert x_{i}\rVert}{\lVert\mathcal{D}\rVert}$ denotes the average number of nonzero entries per vector.
The search process for the full-precision Sindi index is summarized in Algorithm 2, and consists of three stages: product computation, accumulation, and heap update. Given a query $\vec{q}$, inverted index $I$, and target recall size $k$, a distance array $A$ of length $\lambda$ is initialized to zeros (Lines 1–2) and an empty min-heap $H$ is created (Line 3). The outer loop traverses all windows $w\in\{0,\dots,\sigma-1\}$ (Line 4). For each non-zero query component $q^{j}$ (Line 5), SIMD-based batched multiplication is performed with all $x_{i}^{j}$ in $I_{j,w}$, and the results are stored sequentially into a temporary product array $T^{j}$ aligned with $I_{j,w}$ (Line 6). Each $x_{i}^{j}$ is then retrieved (Lines 7–8), its mapped index $m=i\bmod\lambda$ computed (Line 9), and $T^{j}[t]$ accumulated into $A[m]$ (Line 10). After all $q^{j}$ for the current window are processed, the heap update stage begins (Line 12): each $A[m]$ that exceeds the heap minimum or when $H$ has fewer than $k$ entries is inserted into $H$ with its global ID $(m+\lambda w)$ and score $A[m]$ (Lines 13–14), removing the smallest if size exceeds $k$ (Lines 15–16). $A[m]$ is then reset to zero for the next window (Line 17). When all windows are processed, $H$ contains up to $k$ vector IDs with their full-precision distances to $\vec{q}$, which are returned as the final result (Line 20).
Complexity. Let $l=\frac{\sum_{q^{j}\in q}|I_{j}|}{\lVert q\rVert}$ denote the average number of nonzero entries in the traversed lists. With Window Switch, the total number of postings visited is still $\lVert q\rVert\,l$, so the time complexity of a full-precision query is $\mathcal{O}\!\left(\frac{\lVert q\rVert\,l}{s}\right)$, where $s$ is the SIMD batch size. Hence, the computational cost is independent of the window size $\lambda$.
@@ 21    $A[m]\leftarrow 0$
3.4. Analysis of Window Size’s Impact on Performance
@@
  • Random-access cost to the distance array. During accumulation, each partial score $T^{j}[t]$ is added to $A[i\bmod\lambda]$, producing a random write pattern over $A$. When $\lambda$ decreases, the length of $A$ becomes smaller, more of it fits in cache, and this random-access cost decreases due to fewer cache misses.
@@
Figure 5 reports experimental results for the full-precision Sindi on the SPLADE-1M and SPLADE-FULL datasets. For each dataset, we executed 6,9806,\!980 queries under different window sizes $\lambda$ and measured the QPS. We also used the Intel VTune Profiler (Intel Corporation, 2025) to record memory bound metrics for two types of memory accesses: random accesses to the distance array (arising from accumulation writes) and cache eviction to sub-lists (occurring when switching dimensions within a window). Here, memory bound denotes the percentage of execution time stalled due to memory accesses. For the SPLADE-1M dataset, as $\lambda$ increases from 11K to 11M, the memory bound from distance array accesses increases monotonically, whereas that from sub-list cache evictions decreases monotonically. The total memory-bound latency reaches its minimum near $\lambda\approx 100$K, corresponding to the highest query throughput. The SPLADE-FULL dataset exhibits the same trend, confirming the existence of an optimal window size.
Based on this, the memory-access latency for queries can be modeled by a double power-law (Bertsekas and Gallager, 2021; Hennessy and Patterson, 2011):
(2) $T_{\text{mem}}(\lambda)=A\lambda^{+\alpha}+B\lambda^{-\beta}+C,$
@@
  • $A\lambda^{+\alpha}$ models the increasing cost of random accesses to the distance array as $\lambda$ grows;
  • $B\lambda^{-\beta}$ models the decreasing cost of cache eviction to sub-lists with larger $\lambda$;
  • $C$ is the baseline memory-access cost unrelated to $\lambda$.
This function reaches its minimum at $\lambda^{*}=\left(\frac{B\beta}{A\alpha}\right)^{\frac{1}{\alpha+\beta}}$. When $\lambda\ll\lambda^{*}$, $T_{\text{mem}}$ is dominated by the sub-list eviction term and decreases as $\lambda$ increases; when $\lambda\gg\lambda^{*}$, the distance-array term dominates and $T_{\text{mem}}$ grows with $\lambda$. The optimum $\lambda^{*}$ occurs when these two terms are balanced.
@@
Figure 5 shows that the dashed line represents the theoretical QPS curve. It is obtained by estimating $(\alpha,\beta)$ via log–log regression (Draper, 1998) and $(A,B,C)$ via least-squares fitting (Lawson and Hanson, 1995) of the double power-law model. For the SPLADE-1M dataset, the model predicts an optimal $\lambda^{*}\approx 7.35\times 10^{4}$, while for SPLADE-FULL it predicts $\lambda^{*}\approx 1.25\times 10^{5}$. These predictions are reasonably consistent with the measured optimal $\lambda\approx 10^{5}$, with minor discrepancies attributable to hardware-level performance variability and measurement noise introduced by VTune profiling.
() SPLADE-1M
(a) SPLADE-FULL
Figure 5. Impact of Window Size on Query Throughput and Memory Accesses.
4. Approximate Inverted Index
@@
4.1. Pruning Strategies
A key property of sparse vectors is that a small number of high-valued non-zero entries can preserve the majority of the vector’s overall information content (Formal et al., 2021). This distribution pattern typically results from the training objectives of sparse representation models. For instance, SPLADE often concentrates the most informative components into a limited set of high-weight dimensions. Conversely, many low-valued non-zero entries correspond to common stopwords (e.g., “is”, “the”), which can be safely removed with minimal impact on retrieval quality.
Let $\vec{x}_{i}^{\prime}$ denote the pruned version of document $\vec{x}_{i}$, and $\vec{q}^{\prime}$ the pruned version of query $\vec{q}$. Let $l$ be the average posting-list length before pruning and $l^{\prime}$ the average posting–list length after pruning. The reduction in computational cost achieved by pruning is
$\|q\|\cdot l\;-\;\|q^{\prime}\|\cdot l^{\prime},$
For a given pruning operator $\phi$, we define the inner product error for document $\vec{x}_{i}$ as
$e^{(\phi)}_{i}=\delta(\vec{x}_{i},\vec{q})-\delta\!\bigl{(}\phi(\vec{x}_{i}),\phi(\vec{q})\bigr{)},$
where $\delta(\cdot,\cdot)$ denotes the exact inner product and $\phi(\cdot)$ applies the pruning transformation. The total inner product error over the dataset $\mathcal{D}$ is then
$\varepsilon^{(\phi)}=\sum_{\vec{x}_{i}\in\mathcal{D}}e^{(\phi)}_{i}$

### v3

@@ III-B Efficient Distance Computation
Figure 4 illustrates the search procedure. Since $\lVert\mathcal{D}\rVert=9$, the distance array $A$ is initialized with $\text{size}(A)=9$ and all elements set to 0. The query $\vec{q}$ contains three non-zero components, $q^{1}$, $q^{5}$, and $q^{8}$, so only the inverted lists $I_{1}$, $I_{5}$, and $I_{8}$ are traversed. Consider $\vec{x}_{4}$ as an example. From $I_{1}$, we obtain $x_{4}^{1}=6.8$, and the product $q^{1}\times x_{4}^{1}=17.0$ (this multiplication can be SIMD-accelerated) is temporarily stored in $P[0]$. Accumulating $P[0]$ into $A[4]$ gives $A[4]=0+17.0=17.0$. Similarly, from $I_{5}$ we compute $x_{4}^{5}\times q^{5}=14.0$, store it in $P[2]$, and add it to obtain $A[4]=31.0$. The computation for $I_{8}$ proceeds analogously. Finally, $A[4]$ becomes $36.1$, which equals $\delta(\vec{x}_{4},\vec{q})$. Although $\vec{x}_{4}$ has another non-zero entry $x_{4}^{2}$, it is not in $\Omega(\vec{x}_{4},\vec{q})$ and thus does not contribute to the inner product. The same accumulation process applies to $\vec{x}_{2}$, $\vec{x}_{3}$, $\vec{x}_{6}$, and $\vec{x}_{7}$. Eventually, $A[4]$ holds the largest value, so the nearest neighbor of $\vec{q}$ is $\vec{x}_{4}$.
Figure 4: An Example of Sindi and Query Process.
Using SIMD instructions, Sindi processes the $j$-th inverted list $I_{j}$ in batches, multiplying $q^{j}$ with each $x_{i}^{j}$ it contains and writing the results sequentially into $P^{j}$. This approach maximizes CPU utilization and reduces the complexity of the inner product computation from $O(\lVert q\rVert+\lVert x\rVert)$ to $O\!\left(\frac{\lVert q\rVert}{s}\right)$, where $s$ denotes the number of elements processed per SIMD operation. The complexity of Sindi’s distance computation is derived as follows:
Theorem III.1 (Amortized Time Complexity of Sindi Distance Computation).
@@
Proof.
The total number of non-zero entries accessed in all posting lists for $\mathcal{J}$ is $\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert.$ Since $s$ entries can be processed in parallel using SIMD, the total time is $T_{\text{total}}=\sum_{j\in\mathcal{J}}\frac{\lVert\mathcal{I}^{j}\rVert}{s}.$ Amortizing over all $\lVert\mathcal{X}\rVert$ candidates gives $T=\frac{T_{\text{total}}}{\lVert\mathcal{X}\rVert}=\frac{\sum_{j\in\mathcal{J}}\lVert\mathcal{I}^{j}\rVert}{s\cdot\lVert\mathcal{X}\rVert}.$ For any $\vec{x}_{i}\in\mathcal{X}$, we have $q\cap x_{i}\subseteq q$, implying that $\|\Omega(\vec{x}_{i},\vec{q})\|\leq\|q\|$. Therefore: $T\leq\frac{\sum_{\vec{x}_{i}\in\mathcal{X}}\|q\|}{s\cdot\|\mathcal{X}\|}=\frac{\|q\|\cdot\|\mathcal{X}\|}{s\cdot\|\mathcal{X}\|}=\frac{\|q\|}{s}.$ Substituting into $T$ yields $T\leq\frac{\sum_{\vec{x}_{i}\in\mathcal{X}}\lVert q\rVert}{s\cdot\lVert\mathcal{X}\rVert}=\frac{\lVert q\rVert}{s}.$ Hence, the amortized per-vector complexity is $\Theta\!\left(\frac{\lVert q\rVert}{s}\right).$
III-C Cache Optimization
@@ III-C1 Window Switch
(1) Inner product computation. For each scanned posting list $I_{j,w}$, compute the products $q^{j}\times x_{i}^{j}$ for all its entries, using SIMD instructions when possible. These products are written sequentially into a temporary array $P^{j}$, which has the same length as $I_{j,w}$ and is index-aligned with it — i.e., $P^{j}[t]$ stores the product for the $t$-th posting in $I_{j,w}$. The accumulation stage then adds each $P^{j}[t]$ into the corresponding $A[i\bmod\lambda]$ using $O(1)$ time per update.
@@
Note that Window Switch only changes the order of list entries scanned, without altering the number of arithmetic operations. Thus, the time complexity of distance computation remains $O\!\left(\frac{\lVert q\rVert}{s}\right)$.
@@ 4     $w\leftarrow\lfloor\frac{i}{\lambda}\rfloor$
III-C2 Construction and Search
The construction process of the full-precision Sindi index with Window Switch is shown in Algorithm 1. Given a sparse vector dataset $\mathcal{D}$ of dimension $d$, the algorithm iterates over each dimension $j$ (Line 1). For each $j$, it collects all vectors $\vec{x}_{i}\in\mathcal{D}$ having a non-zero entry $x_{i}^{j}$ into a temporary set $\mathcal{X}$ (Line 2). These vectors are then appended to the corresponding window $I_{j,w}$ of $I_{j}$ based on their IDs, where $w=\lfloor i/\lambda\rfloor$ (Lines 3–5). After processing all dimensions, the inverted index $I$ is returned (Line 6). The time complexity of the construction process is $O(\lVert\mathcal{D}\rVert\lVert\bar{x}\rVert)$, where $\lVert\bar{x}\rVert=\frac{\sum_{\vec{x}_{i}\in\mathcal{D}}\lVert x_{i}\rVert}{\lVert\mathcal{D}\rVert}$ denotes the average number of nonzero entries per vector.
The search process for the full-precision Sindi index is summarized in Algorithm 2, and consists of three stages: product computation, accumulation, and heap update. Given a query $\vec{q}$, inverted index $I$, and target recall size $k$, a distance array $A$ of length $\lambda$ is initialized to zeros (Lines 1–2) and an empty min-heap $H$ is created (Line 3). The outer loop traverses all windows $w\in\{0,\dots,\sigma-1\}$ (Line 4). For each non-zero query component $q^{j}$ (Line 5), SIMD-based batched multiplication is performed with all $x_{i}^{j}$ in $I_{j,w}$, and the results are stored sequentially into a temporary product array $P^{j}$ aligned with $I_{j,w}$ (Line 6). Each $x_{i}^{j}$ is then retrieved (Lines 7–8), its mapped index $m=i\bmod\lambda$ computed (Line 9), and $P^{j}[t]$ accumulated into $A[m]$ (Line 10). After all $q^{j}$ for the current window are processed, the heap update stage begins (Line 12): each $A[m]$ that exceeds the heap minimum or when $H$ has fewer than $k$ entries is inserted into $H$ with its global ID $(m+\lambda w)$ and score $A[m]$ (Lines 13–14), removing the smallest if size exceeds $k$ (Lines 15–16). $A[m]$ is then reset to zero for the next window (Line 17). When all windows are processed, $H$ contains up to $k$ vector IDs with their full-precision distances to $\vec{q}$, which are returned as the final result (Line 20).
Complexity. Let $l=\frac{\sum_{q^{j}\in q}|I_{j}|}{\lVert q\rVert}$ denote the average number of nonzero entries in the traversed lists. With Window Switch, the total number of postings visited is still $\lVert q\rVert\,l$, so the time complexity of a full-precision query is $O\!\left(\frac{\lVert q\rVert\,l}{s}\right)$, where $s$ is the SIMD batch size. Hence, the computational cost is independent of the window size $\lambda$.
@@ 21     $A[m]\leftarrow 0$
III-D Analysis of Window Size’s Impact on Performance
@@
  • Random-access cost to the distance array. During accumulation, each partial score $P^{j}[t]$ is added to $A[i\bmod\lambda]$, producing a random write pattern over $A$. When $\lambda$ decreases, the length of $A$ becomes smaller, more of it fits in cache, and this random-access cost decreases due to fewer cache misses.
@@
Figure 5 reports experimental results for the full-precision Sindi on the SPLADE-FULL and AntSparse-10M datasets. For each dataset, we executed queries under different window sizes $\lambda$ and measured the QPS. We also used the Intel VTune Profiler [vtune] to record memory bound metrics for two types of memory accesses: random accesses to the distance array (arising from accumulation writes) and cache eviction to sub-lists (occurring when switching dimensions within a window). Here, memory bound denotes the percentage of execution time stalled due to memory accesses. For the SPLADE-FULL dataset, as $\lambda$ increases from 11K to 8.88.8M, the memory bound from distance array accesses increases monotonically, whereas that from sub-list cache evictions decreases monotonically. The total memory-bound latency reaches its minimum near $\lambda\approx 150$K, corresponding to the highest query throughput. The AntSparse-10M dataset exhibits the same trend, confirming the existence of an optimal window size.
Based on this, the memory-access latency for queries can be modeled by a double power-law [bertsekas2021data, hennessy2011computer]:
$T_{\text{mem}}(\lambda)=c_{rand}\lambda^{+\alpha}+c_{evict}\lambda^{-\beta}+C_{0},$ (1)
@@
  • $c_{rand}\lambda^{+\alpha}$ models the increasing cost of random accesses to the distance array as $\lambda$ grows;
  • $c_{evict}\lambda^{-\beta}$ models the decreasing cost of cache eviction to sub-lists with larger $\lambda$;
  • $C_{0}$ is the baseline memory-access cost unrelated to $\lambda$.
This function reaches its minimum at $\lambda^{*}=\left(\frac{c_{evict}\beta}{c_{rand}\alpha}\right)^{\frac{1}{\alpha+\beta}}$. When $\lambda\ll\lambda^{*}$, $T_{\text{mem}}$ is dominated by the sub-list eviction term and decreases as $\lambda$ increases; when $\lambda\gg\lambda^{*}$, the distance-array term dominates and $T_{\text{mem}}$ grows with $\lambda$. The optimum $\lambda^{*}$ occurs when these two terms are balanced.
@@
Figure 5 presents the theoretical QPS curves, where parameters $(\alpha,\beta)$ are estimated via log–log regression [draper1998applied] and $(c_{rand},c_{evict},C_{0})$ via least-squares fitting [lawson1995solving] of the double power-law model. For SPLADE-FULL, the model predicts an optimal $\lambda^{*}\approx 1.2\times 10^{5}$, while for AntSparse-10M, it predicts $\lambda^{*}\approx 5.1\times 10^{4}$. Both predictions align with the robust high-throughput interval $\lambda\in[50,000,100,000]$. This broad plateau confirms that within this range, performance is stable, rendering precise point-wise tuning unnecessary despite minor hardware-level variability. More details are shown in Appendix B.
() SPLADE-FULL
(a) AntSparse-10M
Figure 5: Impact of Window Size on Query Throughput and Memory Accesses.
IV Approximate Inverted Index
@@
IV-A Pruning Strategies
A key property of sparse vectors is that a small number of high-valued non-zero entries can preserve the majority of the vector’s overall information content [splade]. This distribution pattern typically results from the training objectives of sparse representation models. For instance, SPLADE often concentrates the most informative components into a limited set of high-weight dimensions. Conversely, many low-valued non-zero entries correspond to common stopwords (e.g., “is”, “the”), which can be safely removed with minimal impact on retrieval quality.
Let $\vec{x}_{i}^{\prime}$ denote the pruned version of document $\vec{x}_{i}$, and $\vec{q}^{\prime}$ the pruned version of query $\vec{q}$. Let $l$ be the average posting-list length before pruning and $l^{\prime}$ the average posting–list length after pruning. The reduction in computational cost achieved by pruning is $\|q\|\cdot l\;-\;\|q^{\prime}\|\cdot l^{\prime},$
For a given pruning operator $\phi$, we define the inner product error for document $\vec{x}_{i}$ as $e^{(\phi)}_{i}=\delta(\vec{x}_{i},\vec{q})-\delta\!\bigl(\phi(\vec{x}_{i}),\phi(\vec{q})\bigr),$ where $\delta(\cdot,\cdot)$ denotes the exact inner product and $\phi(\cdot)$ applies the pruning transformation. The total inner product error over the dataset $\mathcal{D}$ is then $\varepsilon^{(\phi)}=\sum_{\vec{x}_{i}\in\mathcal{D}}e^{(\phi)}_{i}$